OpenStudy (anonymous):

Evaluate: 5 • 4 − 8 ÷ 2

OpenStudy (anonymous):

Assuming no parentheses, we have to start by doing multiplication and division, in order from left to right. So after multiplying the first two values: 20 - 8 / 2 Now you need to divide, can you take it from here?

OpenStudy (anonymous):

8/2 =4

OpenStudy (anonymous):

16?

OpenStudy (anonymous):

Start with 5 *4 what you get ?

OpenStudy (anonymous):

20 right ? then 8 / 2 = 4 right ?

OpenStudy (anonymous):

5*2=20

OpenStudy (anonymous):

so 20 -4 = ?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

BAM ! that your answer !

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

yw :D